OpenStudy (anonymous):

A parallelogram has an area of 42 cm2. What would the area be if the base was one-third as long and the height was twice as long?

OpenStudy (apoorvk):

area of a parallelogram 'A' = base 'b' x height'h' or, A=bh now new base =b/3 new height =2h therefore new area = 2h x b/3 =(2/3)bh = two-thirds of the original area.

OpenStudy (campbell_st):

base x height = 42 1/3 base x 2 height = 1/3 x 2 x (base x height) = 2/3(base x height) = 2/3 x 42 Area = 28

OpenStudy (apoorvk):

=(2/3) of 42 , that is.